OpenStudy (anonymous):

what are the steps to finding (5m^2-4p^2)^2

OpenStudy (anonymous):

(5m^2-4p^2)^2 =(5m^2-4p^2)(5m^2-4p^2) =25m^4-20mp^2-20mp^2+16p^4 =25m^4-40mp^2+16p^4

OpenStudy (campbell_st):

this is binomial expansion (5m^2 - 4p^2)(5m^2 - 4p^2)